OpenStudy (anonymous):

Use scientific notation to simplify the expression. Give the answer in standard notation 288,000 ------------ (12,000)(0.00004)

OpenStudy (anonymous):

\[= (2.88\times 10^5)/ ((1.2 \times 10^4)(4.0 \times 10^{-5}))\] \[= ((2.88\times 10^5)(4.0 \times 10^{5}))/(1.2 \times 10^{4})\] \[= ((2.88)(4.0)/1.2) \times 10^{6}\] \[= 0.6 \times 10^{6}\] \[= 6.0 \times 10^{5}\]